A Go-ped traveled 9 miles in 24 minutes.
How long would it take to go 27 miles?
A.) 1 hour, 6 minutes
B.) 1 hour, 12 minutes
C.) 1 hour, 26 minutes
D.) 1 hour

Answer by Nate(3500)

Distance = Rate * Time
Rate = Distance / Time
Rate = 9 / 24 = 3 / 8
Distance = Rate * Time
Time = Distance / Rate
Time = 27 / (3 / 8) = 27(8 / 3) = 9 * 8 = 72 minutes
1 hour and 12 minutes
B.)